Deck A has 30 red cards and 22 black cards. Deck B has 26 red cards and 26 black cards. A card is drawn at random from Deck A and transferred to Deck B. Now, a card is drawn at random from Deck B.
Find the probability that the transferred card from Deck A is red and the drawn card from deck b is also red after card transfer.

Answer by ewatrrr(24785)


Hi  
Deck A has 30 red cards and 22 black cards: 52 cards 
Deck B has 26 red cards and 26 black cards: 52 cards

' A' card is drawn at random from Deck A 
(and transferred to Deck B)
P(red) = 30/52 = 15/26

Deck B NOW has 27 red cards and 26 black cards: 53 cards
'a' card is drawn at random from Deck B.
P(red) = 27/53 
Wish You the Best in your Studies.